OpenStudy (anonymous):

Claculate the percent markdown for a product whose original price was $249.99 marked down to $149.99

OpenStudy (anonymous):

umm this is eng but... 249.99 X Y=149.99 Y=149.99/249.99 solve Y

OpenStudy (anonymous):

y=mx+b ? is that what your trying to do ???

OpenStudy (anonymous):

your looking for houw much you save so solve for how much is 149 of 249 than subtract that number from 100

OpenStudy (anonymous):

y=mx+b is for straight line in a graph. what your asking is simple math but you should really be in the math section

OpenStudy (lanre):

You may not get accurate answer here. Please post this question in the mathematics section. Click on the Find More Subject at the top, scroll down to select mathematics, click to navigate to the page.

OpenStudy (anonymous):

y=mx+b is not always a strait line formula for a graph.

OpenStudy (anonymous):

simple math can be messed up with a crazy formula that is not needed .

OpenStudy (anonymous):

100.00 = X% * 249.99

OpenStudy (anonymous):

So... 100.00/249.99 = X% * 249.99/249.99 Then... 40.0 = X%

OpenStudy (trancenova):

Hello, this question is in the wrong section, so it's not close. Please ask questions in the correct section, find more sections by clicking "find more subjects" above. If you have any questions feel free to ask :)